# It Is Time To Steer: A Scalable Framework for Analysis-driven Attack Graph Generation

### v1

@@ 6 Case Study Evaluation
In this section, we illustrate a case study to show the capabilities of the approach to be applied in a real setting. We first show how the approach can be used in real-time with a network bigger than the ones typically analyzed by state-of-the-art solutions allowing higher scalability. Next, we provide a systematization of attack path analyses retrieved from the literature and show how the approach is capable of supporting them.
6.1 Application to large networks
The first application we prove deals with the scalability of the networks. In Table I we report the works of the literature that address the generation and analysis of attack paths, highlighting their scalability in terms of the network parameters. In particular, we consider the maximum number of hosts, vulnerability per host, and whether they generate all paths or not. In this last case, the approaches define heuristics to prioritize the paths under analysis and avoid the full generation (more details in Section 7).
TABLE I: Scalability parameters of existing works on attack path computation. ND=Not Defined
@@ 6.1 Application to large networks
From Table I we can observe that the largest experimental settings include either many hosts (50) and few vulnerabilities (5), or a medium number of hosts (15) and vulnerabilities (12 vulnerabilities per host). For the validation of our approach, we already considered networks bigger than those (see Section 4.1.1), but here we want to report an even bigger case study with 100 hosts and 100 vulnerabilities per host, all different from each other to prove our approach provides a big improvement over the state-of-the-art. Concerning the attack path query and for the sake of example, we consider the most common risk analysis that requires attack paths with the highest risk, therefore with both likelihood and impact between 0.9 and 1, as the most risky situations. We prove later the support of a wide range of analysis queries. Let us note that the generation of the complete AG is intractable for such a size with the classic full generation, but we show that the progressive generation allows having partial results even in this case, through the real-time indicators that are the stability (for StatAG) and precision (for SteerAG) controllable in real-time by an analyst or an automatic process.
Figure 9: Attack Path features stability trend during the execution of StatAG.
@@
From the stability trend of Fig. 9 it is evident that until iteration 100 the results are early partial, with very low and fluctuant stability indicating a still high approximation of the complete AG. The analyst can no matter what start exploring in this phase, even by iteration 1, and eventually query the system for her information needs. Contrary, from iteration 100 s/he starts observing high stability values both for likelihood and impact, thus can hypothesize that it is the time of significant partial results, and support tasks like hypotheses formulation. In a real scenario, it is reasonable to observe a constant trend of the stability values before considering the results statistically significant. From the observation of Fig. 9, a reasonable time to consider mature partial results is at iteration 300, where both likelihood and impact show a regular median stability of 0.85, progressively increasing. This phase supports hypothesis testing, comparative or confirmatory analyses, and early decision-making. From this time, the attack path analysis can start with the consideration that it approximates at least 85% of the complete AG.
Figure 10: Precision trend during the execution of SteerAG.
@@
In the meanwhile, the analyst observes the precision trend of Fig. 10. S/he notices a rapid growth at iteration 100, indicating that the SteerAG has been activated. This observation is in line with the stability trend (Fig. 9), thus indicating the starting point of mature partial results. The precision has an average value of 0.8 until around iteration 8300. During this long period, the analyst can perform attack path analysis, considering that new paths are incoming to answer the query and, therefore, s/he should progressively monitor the new results. When the precision is stable for a longer period (i.e., around iteration 2000), s/he can consider that most of the attack paths are included in the query and, from this moment on, assumes that the analysis is composed of definitive partial results (i.e., other incoming paths do not significantly change the attack path feature distributions. Finally, when the precision breaks down around iteration 8300, the results are very close to the exact ones. Let us note, that this observation agrees with the stability trend (Fig. 9), indicating from iteration 1000, a stability very close to 1, until the end of the execution. From this iteration on, the analyst can produce the final outcomes of the analysis and proceed with her work.
To give an idea of the advantages provided by the proposed framework, we identify two potential targets: a human analyst and an automatic analysis process. The former takes advantage of the progressive generation to perform early decision-making in interactive times. In fact, each iteration has an average duration between 0.5 and 5 seconds, and after a few minutes ($\approx 5$) from the beginning of execution, s/he can start performing analysis with mature partial results. The latter has the advantage of progressive results feeding, without the need for interactive response time but instead based on results accuracy thresholds: the attack path analysis reaches definitive results after a few hours ($\approx 12$) making it computable in contrast to classic approaches that do not support it.
@@ 6.2 Coverage of Attack Path Analyses
To conclude the presentation of the case study, we analyze how the proposed approach can analyze a selected set of queries with the configuration settings used for the validation (Section 4.1.1). The attack path queries are extracted from state-of-the-art works addressing the systematization of attack path analysis [57, 58], and they are the following:
@@
Q5: Retrieve the attack paths corresponding to black swan attacks, i.e., those ones with very high impacts (impact $>$ 0.9), but very unlikely to happen (likelihood $<$ 0.3) [62];
Q6: Retrieve the attack paths corresponding to gray swan attacks, i.e., those ones with very low impacts (impact $<$ 0.3), but highly probable (likelihood $>$ 0.9) [63];
Q7: Prioritize the attack paths by risk [8]. Let us note that this query corresponds to an ensemble of queries, retrieving first the paths with risk 1, then between 0.9 and 1, and so on. To avoid the complete enumeration of the paths, we stop the approach once we retrieve the paths with risk values higher than 0.5.
Fig. 11 reports the recall median trend for the analysis of these queries, where each experiment is repeated 100 times.
Figure 11: Recall trend for the attack path analyses of the query systematization.
The recall trends in Fig. 11 show that most of the queries (specifically, Q1-Q5 and Q7) converge very rapidly to the GT, indicating the concrete support that the approach can provide during attack path analysis. Among them, Q2, Q3, and Q4 ask for precise values of attack path features, while Q1, Q5, and Q7 ask for more complex queries, requiring more than one traversal in the same AG. In particular query Q6 for gray swan attacks has the worst performance requiring 3500 median iterations to retrieve all the paths. The reason for its worse performance may be the need to perform more traversals to answer the query, which includes many paths. Let us remark that it still provides approximate answers but with a slower convergence rate.
This case study showed the capabilities of the approach to perform attack path analysis with networks’ size that existing solutions cannot address and according to state-of-the-art attack path analysis.
@@ 7 Related Work
Several works leverage Attack Graphs (AGs) for analyzing cyber risks of computer networks [8, 9, 64], which underline their expressiveness and high computational complexity. Our framework intersects two main areas of the related work: AG generation and attack path analysis.
7.1 Attack Graph Generation
To address the scalability issue of AGs, the research focuses on the design of approaches to generate AGs.
@@ 7.1 Attack Graph Generation
One possible solution consists of leveraging distributed and parallel computing. The distributed approaches, as Kaynar et al. [11] and Sabur et al. [65], partition the network based on its services and vulnerabilities and assign the different partitions among distributed agents. In this way, each agent computes the attack paths on a smaller portion of the graph which is finally combined with the other paths to complete the generation. The approaches based on parallel computing [16, 56] instead adapt the graph search algorithms (i.e., BFS and DFS), that are designed for serial computation, to parallel algorithms that allow running the AG generation in multi-core environments. Our approach is agnostic to potential distribution or parallelization, as it acts on the workflow of analysis and generation.
@@
Another possibility is using Artificial Intelligence (AI). For example, Li et al. [66] leverage Deep Learning and node2vec [67] to train a neural network with data from system logs and use it to predict dependencies between network hosts and label the sequence of events as potential attack paths. Besides, Ychao et al. [15] model the path discovery problem as a planning problem on graphs. They translate vulnerability information to formalism actions and use state-of-the-art planners to discover attack paths.
@@
While Distributed and AI-based solutions provide a good improvement in the scalability of AG generation, they still require waiting until the generation is completed to perform the first analyses. In dynamic scenarios where network components change rapidly (and so do attack graphs), it is not reasonable to wait for the re-computation of the attack paths each time the environment changes.
7.2 Attack Path Analysis
To make AGs actionable for cyber risk analysis, different works addressed the scalability problem from the attack path analysis perspective. The central idea behind these works is the design of heuristics that prioritize the analysis of certain attack paths according to specific security requirements.
@@ 7.2 Attack Path Analysis
Some of them prioritize the attack path analysis considering the graph topological structure: it is the case of Sun et al. [14], who use nodes’ in-degree and out-degree, and Gonda et al. [68, 34], who map AGs to planning graphs and leverages the nodes’ centrality. Other works define priorities according to security metrics, as Liu et al. [69] who map potential exploits to the easiest reachable attack target, and then prioritize the analysis based on their probability. Similarly, Wang et al. [53] prioritize attack paths on the degree of matching between attack nodes and a set of predefined security conditions, while Feng et al. [12] associate a cost value to each attack node and prioritizes the attack paths based on the costs. These works take specific assumptions to assign priorities, and therefore they are fine only for scenarios that conform to such assumptions. Moreover, most of them are usually static assumptions, which do not necessarily follow the network exposure.
@@
Few works addressed the problem of considering attack path analysis to prioritize the generation. One of them is proposed by Yuan et al. [27] and Salayma et al. [70]. They model the AG through a graph database and perform attack path analysis by suitably writing queries to the database. While these solutions allow the expression of complex and highly customizable queries, they have the drawback that depend on the graph database, which may slow down the performance during the analysis [71]. Nadeem et al. [72] use alerts coming from Intrusion Detection Systems (IDS) to drive the attack path analysis. They translate alert events to episode sequences that are then used to build the AG and execute the analysis driven by alerts. Similarly, Hassan et al. [73] perform rule matching on system logs to identify the events that match attack behaviors described in security knowledge bases. While these alert-based approaches are context-aware, they may not represent the entire environment, such as network components with no alerts detected.
8 Limitations and Opportunities
This paper proposed a new framework to generate and analyze AGs based on progressive data analysis. To the best of the authors’ knowledge, it is the first contribution to advancing the classic sequential AG-based process.
In this section, we elicit some limitations as well as opportunities for the proposed work.
First, we point out that our approach and its validation are applied to topological attack graphs [12, 20, 74], where nodes represent the attacker states on the network and edges the lateral movements of vulnerability exploits. However, there exist logical AGs [22, 75, 31], that model the attacker states according to the dependency relations expressed as logical axioms and formal rules. We believe that our approach can be easily adjusted to perform attack path analysis over logical AGs, for example, adapting the attack path construction with model-checking [76]. We consider this extension as future work.
@@ 8 Limitations and Opportunities
Additionally, we considered the likelihood and impact as attack path features since they are the main ones used in the existing literature on cyber risk management. Nonetheless, there exist different risk models that use different metrics, especially in domains such as cloud computing [77] and automotive [78]. These metrics may present a non-linear relation to the vulnerability features, thus making the decision tree training more challenging. To this aim, we suggest exploring other ML models, in particular Graph Neural Networks [79], that are challenging to adapt to our approach due to the large amount of needed resources for training, the long training time not compatible with interactive analysis, and their black-box nature that make it difficult to retrieve the steering rules.
Finally, another point worth discussing is the possible bias introduced by SteerAG. In fact, the steering rules tend to bias the distribution of vulnerability and attack path features towards the query requirements, as we expected since it actually accelerates the convergence to the complete AG. However, it would be interesting to study a way to keep the two distributions distinct, allowing both to answer correctly the analysis query and at the same time extract from its answer the part to add to statAG to keep its statistical significance and minimize introduced bias.
Other considerations related to the proposed approach examine the opportunities it opens for further research. One involves the improvement of the decision tree predictions after the precision breakdown. Currently, we just re-train the decision tree with the newly incoming paths, so as to update the steering rules. However, more advanced approaches can be studied to detect the precision breakdown and efficiently re-train the ML model. A solution that we want to explore for this aim is Network Representation Learning [80].
An additional opportunity is to study the smart combination of analysis queries to create strategies for the generation of the AG, prioritizing the analyst goal by gaining a higher versatility than classic sampling methods like stratified sampling. These combinations can be initially executed even without the security operator’s intervention, obtaining a highly dynamic way of prioritizing the AG to generate.
Finally, while the proposed framework can be included in an automatic pipeline, where a process issues analysis queries, it opens the research for the design of human-in-the-loop systems to support security experts during the real-time risk analysis of attack paths, possibly with some form of active learning [81, 82] where human interactions help accelerate the AG generation and analysis.
9 Conclusion
This paper presented a new framework for AG-based systems addressing the problem of their scalability by leveraging progressive data analysis. We designed two approaches to progressively generate statistically significant partial AGs (StatAG) and accelerate the generation based on the attack path query (SteerAG). Additionally, it can better adapt to network changes (e.g., new connections or vulnerabilities in the network) without the need for recomputing the entire AG, a benefit that classic AG generation approaches lack. We provided extensive experimental validation and showed a case study demonstrating the capability of the framework to (i) analyze networks whose size cannot be managed with existing solutions, and (ii) efficiently support common attack path analyses. The approach is orthogonal to the existing ones and can be used in their combination without loss of generality.
In future works, we intend to i) Investigate the application of StatAG and SteerAG to the class of logical AGs, ii) extend it to more heterogeneous risk models, iii) apply advanced learning solutions, such as network representation learning, and iv) develop an interactive system to validate the approach in real scenarios with security experts.
Appendix A Attack Path Sampling Analysis
(a) BFS.
(b) DFS.
(c) Random Walk.
Figure 12: KS distance trend of partial AGs from the GT for the attack path features distribution when sampling using (a) BFS, (b) DFS, and (c) Random Walks algorithms.
In this appendix, we report the analysis that shows the better performance of random walk sampling than BFS and DFS ones. We use the experimental setting defined in Section 4.1.1 and report in Fig. 12 the comparison of the convergence trend of the KS distance of partial AGs from the complete AG (i.e., the ground truth GT) for the three sampling algorithms.
Considering the different trends, we can observe that the BFS sampling (Fig. 12a) has the worst performance for two main reasons. First, the results of the experiments have higher variability if compared with DFS 12b and random walk 12c, indicating less controlled KS distances from the GT. Second, the median values have a higher constant trend of 0.2 that increases further for the impact. In contrast, random walk sampling has better performance, with low data variability right from the first iterations and with monotonically decreasing convergence. Finally, DFS sampling has lower variability than BFS, but the median values are at 0.2, while random walks represent well the complete AG with a median KS distance closer to 0.
This analysis is in line with our expectation, because both BFS and DFS sampling are biased to the starting node, and as a consequence, they tend to visit the same edges and, therefore, the same vulnerabilities. Contrary, random walks enable visiting the graph in an unbiased manner, thus increasing the unique vulnerability visited and this results in a faster convergence to the attack path features of the complete AG. For this reason, we included random walk sampling in our approach.
Appendix B An example of Attack Graph
In this appendix, we provide an example of an attack graph, modeled according to the AG model described in section 2.
(a) Simple network configuration.
(b) Attack Graph for the simple network configuration.
Figure 13: Example of (a) network configuration and (b) its resulting attack graph model.
For the sake of explanation, let us consider the simple network reported in Fig. 13a [74], in which the Demilitarized Zone (DMZ) is safeguarded by two firewalls, namely Firewall-1 and Firewall-2. It has a Web Server (H1) and a Login Server (H2). Firewall-1 shields the DMZ from external threats originating from the attacker located on Host H0, and it allows only http and ssh traffic. Firewall-2 safeguards the DMZ from internal threats and permits access to the Database server solely from the Web server. The vulnerability inventory is reported in Table II, while the resulting attack graph is in Fig. 13b.
TABLE II: Vulnerability inventory for the illustrative example.
Software Host Vulnerability
Apache Web Server H1 V1 (CVE-2006-3747)
OpenSSH H2 V2 (CVE-2002-0640)
MySQL H3 V3 (CVE-2009-2446)
Linux Kernel H3 V4 (CVE-2004-0495)
Linux Kernel H3 V5 (CVE-2023-3269)
Host H1 is running a vulnerable version of Apache web server, which has a vulnerability (CVE-2006-3747) that allows a remote attacker to exploit and gain user privilege on the Web Server. The ssh service on H2 has a vulnerability (CVE-2002-0640), which allows remote attackers to gain user privilege. Database server H3 is a Linux box running MySQL database which has a remotely exploitable vulnerability (CVE-2009-2446), enabling the attacker to gain user privilege. The Linux kernel in host H3 also has a vulnerability (CVE-2004-0495) that allows local users to gain root privilege.
Let us note that vulnerabilities V4 and V5 affect the same service, resulting in multi-edges in the attack graph.
Appendix C Risk Model details
This appendix provides additional details of the risk model introduced in Section 2. It considers CVSS standard metrics to estimate the likelihood and impact of an attack path. In particular, the likelihood is calculated according to the following formulas:
$exploitability(u)=RoundUp(AV\cdot AC\cdot PR\cdot UI)$ (4)
$likelihood(AP)=norm(\sum_{u\in U_{AP}}\frac{1}{exploitability(u)})$ (5)
Equation 4 determines the exploitability of a vulnerability $u$ by multiplying the CVSS-3.1 exploitability metrics, that are :
  • Attack Vector (AV): measure the context by which vulnerability exploitation is possible, thus a larger value corresponds to a more remote attacker;
  • Attack Complexity (AC): measures the conditions beyond the attacker’s control that must exist to exploit a vulnerability;
  • Privilege Required (PR): measures the level of privileges an attacker must possess before successfully exploiting a vulnerability;
  • User Interaction (UI): measures the requirement for a human attacker to successfully compromise a vulnerable component.
Given that an attack path $AP$ is composed of multiple vulnerabilities, then its likelihood is calculated as the normalized sum of the inverse of exploitability, as reported in Equation 5. This choice is due to the CVSS exploitability metrics, defined in the range [0,1], which have the rationale that the easier the exploitation of the vulnerability, the closer they are to one. Thus the likelihood is measured as the inverse of their product, summed for all the vulnerabilities in the path, normalized to be in the range [0,1].
The impact of an attack is determined by the last node in the path, as it represents the attacker’s final goal. In particular, let $u_{n}$ be the last vulnerability of an attack path $AP$ which allows it to finally reach the attacker target. Then,
$impact(AP)=1-[(1-C_{n})\cdot(1-I_{n})\cdot(1-A_{n})]$ (6)
$risk(AP)=likelihood(AP)\cdot impact(AP)$ (7)
where the CVSS-3.1 impact metrics are used, that are:
  • Confidentiality (C): measures the impact on information access and disclosure.
  • Integrity (I): measures the impact on trustworthiness and veracity of information;
  • Availability (A): measures the impact on accessibility of information resources.
Contrary to exploitability metrics, the CVSS impact metrics have the rationale that the more damages caused by successful exploitation, the closer they are to 1. For this reason, the impact is calculated by subtracting the value of the metrics from 1. Finally, equation 7 expresses the risk of an attack path $AP$ as the dot product of its likelihood and impact [29].

### v2

@@ 6 Case Study Evaluation
6.1 Application to Large Real Networks
Table 1: SoA settings.
@@ 6.1 Application to Large Real Networks
Proposed approach >250 >50
The first application we prove deals with scalability. In Table 1 we report the works of the literature that address the generation and analysis of attack paths, highlighting the considered maximum number of hosts, vulnerability per host, and whether they generate all paths. In this last case, the approaches define heuristics to prioritize the paths under analysis and avoid the full generation (more details in Section 7). From Table 1 we can observe that the largest experimental settings include either many hosts (50) and few vulnerabilities (5), or a medium number of hosts (15) and vulnerabilities (12 vulnerabilities per host). For the validation of our approach, we already considered networks bigger than those, but here we report an even bigger and real network. We emulated the network of a university department composed of around $250$ devices arranged in $5$ LANs connected through routers. The network has been scanned for vulnerabilities using commercially available vulnerability scanners (i.e., Tenable Nessus).
Figure 8: AP features stability trend.
@@
Concerning the attack path query and for the sake of example, we consider the most common risk analysis that requires attack paths with likelihood and impact higher than 0.9, as the most risky situation. We prove later the support of a wide range of analysis queries. Let us note that the generation of the complete AG is intractable for such a size with the classic full generation, but we show that the progressive generation allows having controllable partial results. The stability trend of Fig. 8 shows that until iteration 100 the results are early partial, with low and fluctuant stability indicating a still high approximation of the complete AG. From iteration 100, high stability values start both for likelihood and impact, thus hypothesizing that it is the time of significant partial results, and support tasks like hypotheses formulation. However, it is reasonable to observe a constant trend of the stability values before considering the results statistically significant: a reasonable time to consider mature partial results is at iteration 500, where both likelihood and impact show a regular median stability of 0.85, progressively increasing. This phase supports hypothesis testing, comparative or confirmatory analyses, and early decision-making. From this time, the attack path analysis can start, considering that it approximates at least 85% of the complete AG, until reaching very stable results (definitive partial) at iteration 8000.
Figure 9: Precision trend SteerAG.
@@
Meanwhile, the analyst observes the precision trend of Fig. 9. She notices a rapid growth at iteration 100, indicating that the SteerAG has been activated, thus indicating the starting point of mature partial results. The precision has an average value of 0.8 until around iteration 8000. During this period, the analyst can perform attack path analysis, considering that new paths are incoming to answer the query. When the precision breaks down around iteration 8000, the results are very close to the exact ones (definitive partial results). These observations agree with the stability trend (Fig. 8). To give an idea of the advantages provided by this framework in terms of time, an analyst uses the progressive generation to perform early decision-making in interactive times. Each iteration has an average duration between 0.5 and 5 seconds, and after a few minutes ($\approx 5$) from the beginning, she can start the analysis with mature partial results for any issued query.
@@ 6.2 Coverage of Attack Path Analyses
To conclude the case study, we analyze how the proposed approach can analyze a selected set of queries with the configuration settings used for the validation (Section 4.1). The attack path queries are extracted from state-of-the-art works addressing the systematization of attack path analysis [22, 38]:
@@
Q5: Retrieve the attack paths corresponding to black swan attacks, i.e., those ones with high impacts (impact $>$ 0.9), but unlikely to happen (likelihood $<$ 0.3) [2];
Q6: Retrieve the attack paths corresponding to gray swan attacks, i.e., those with very low impacts (impact $<$ 0.3), but highly probable (likelihood $>$ 0.9) [19];
Q7: Prioritize the attack paths by risk [49]. Let us note that it corresponds to an ensemble of queries, retrieving first the paths with risk 1, then between 0.9 and 1, and so on. To avoid the complete enumeration of the paths, we stop the retrieval at risk 0.5. Fig. 10 reports the recall median trend for the analysis of these queries, where each experiment is repeated 100 times.
Figure 10: Recall for different queries.
The recall trends in Fig. 10 show that most queries (specifically, Q1-Q5 and Q7) converge very rapidly to the GT, indicating the concrete support the approach can provide during attack path analysis. Among them, Q2, Q3, and Q4 ask for precise values of attack path features, while Q1, Q5, and Q7 ask for more complex queries, requiring more than one traversal in the same AG. In particular, query Q6 for gray swan attacks has the worst performance, requiring 3500 median iterations to retrieve all the paths. The reason for its worse performance is the need to perform more traversals and the many paths answering the query. Let us remark it still provides approximate answers but with slower convergence. This case study showed the capabilities of the approach to perform attack path analysis with networks’ size that existing solutions cannot address and according to state-of-the-art attack path analysis. As a final consideration, let us remark that the queries can be issued anytime during the execution and the approach triggers the steering mechanism upon the arrival of a query.
@@ 7 Related Work
Attack Graph Generation. To address the scalability issue of AGs, one possible solution consists of leveraging distributed and parallel computing. The distributed approaches, as Kaynar et al. [18] and Sabur et al. [36], partition the network based on its services and vulnerabilities and assign the different partitions among distributed agents. Each agent then computes the attack paths on smaller portions of the AG which are finally combined. The approaches based on parallel computing [23, 24] adapt the serial graph search algorithms to multi-core environments. Our approach is agnostic to distribution or parallelization, as it acts on the AG workflow.
@@
Another possibility is using Artificial Intelligence (AI). For example, Li et al. [26] leverage Deep Learning and node2vec [10] to train a neural network with data from system logs and use it to predict dependencies between network hosts and label the sequence of events as potential attack paths. Besides, Ychao et al. [46] model the path discovery problem as a planning problem on graphs. They translate vulnerabilities into formal actions and use state-of-the-art planners to discover attack paths.
@@
While Distributed and AI-based solutions provide a good improvement in the scalability of AG generation, they still require waiting until the generation is completed to perform the first analyses. In dynamic scenarios where network components change rapidly (and so do attack graphs), it is not reasonable to wait for the re-computation of the attack paths each time the environment changes. In contrast, the proposed approach allows for querying Attack Path analyses at any time with quality-controlled statistically significant results.
@@
Attack Path Analysis. Most current AG-based systems have tackled the scalability issue in cyber risk analysis by prioritizing specific attack paths based on security needs. Some of them prioritize the analysis considering the graph’s topological structure, as Sun et al. [41] who use nodes’ degree, and Gonda et al. [8] who map AGs to planning graphs and leverage the nodes’ centrality. Other works leverage security metrics to map potential exploits to the easiest reachable attack target [28] or to prioritize attack paths on the security conditions [6] and their cost values [6]. These works take specific assumptions on priorities and, therefore, are fine only for scenarios that conform to them. Moreover, they are static and do not follow the network exposure.
@@
Few works addressed the problem of considering attack path analysis to prioritize the generation. Yuan et al. [47] and Salayma et al. [37] model the AG through a graph database and perform attack path analysis by suitably writing queries to the database. While these solutions allow the expression of complex customizable queries, they have the drawback that they depend on the graph database, which may slow down the performance during the analysis [11]. Differently, Nadeem et al. [31] use intrusion alerts to drive the attack path analysis. They translate alert events to episode sequences used to build the AG and execute the analysis driven by alerts. Similarly, Hassan et al. [12] perform rule matching on system logs to identify the events that match attack behaviors described in security knowledge bases. While these alert-based approaches are context-aware, they may not represent the entire environment, such as network components with no alerts detected. In contrast, the proposed approach assesses the entire network without introducing any bias to a fixed analysis. On the contrary, it dynamically adapts to different queries.
8 Discussion and Concluding Remarks
This paper proposed a new framework to generate and analyze AGs based on progressive data analysis. We developed StatAG to generate statistically significant partial AGs and SteerAG to accelerate AG generation based on attack path queries. Our approach adapts better to network changes without recomputing the entire AG, a feature lacking in traditional AG generation methods. Extensive experiments and a case study validate its capability to analyze large networks and efficiently support common attack path analyses. While our approach is complementary to existing methods that can be integrated without loss of generality, to the best of the authors’ knowledge, it is the first contribution to advancing the classic AG-based process. All the materials and the source code of all components are publicly available [^4] https://github.com/XAIber-lab/ProgressiveAttackGraph.
@@ 8 Discussion and Concluding Remarks
Some implications of this work are worth discussing. Although our approach is applied to topological AGs, we believe it can be easily adjusted for attack path analysis over logical ones, for example, adapting the path construction with model-checking [3]. Additionally, we considered the likelihood and impact as attack path features since they are the main ones used in the literature on cyber risk management. Still, others could be considered, dependent on the application domain (e.g., cloud [34] and automotive [29]). We also suggest exploring other ML models, e.g., Graph Neural Networks [45] that, while allowing to capture non-linear relations, are challenging to adapt to our approach due to their resource-intensive training and their black-box nature that makes it difficult to retrieve the steering rules.
Looking at the research opportunities the proposed work opens, one concerns the management of progressive thresholds to indicate early, mature, and definitive partial results. Future works may study automated ways to set progressive thresholds based on real-time indicators. Another opportunity involves the improvement of the decision tree predictions after the precision breakdown. Currently, we re-train the decision tree with the newly incoming paths, but more advanced strategies to optimize the re-train of the ML model may be adopted. Finally, while the proposed framework can be included in an automatic pipeline, where a process issues analysis queries, it opens the research for the design of human-in-the-loop systems.
Acknowledgements.
This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
Appendix 0.A Query stringency analysis
We are interested in studying how a stricter query affects performance to avoid situations where a very specific query gets stuck due to slow convergence. To validate this, we define three ranges of queries: low, when the values of the attack path features are in a range of at most 0.2 (strict query); medium, when it is between 0.3 and 0.5; and high, when it is over 0.5 (large query). We report different analyses in Fig. 11.
(a) Recall trend.
(b) Convergence speed.
(c) Missing rate.
Figure 11: Analysis of steering mechanism for different query ranges.
Fig. 11a shows the recall trend of the different experiments for each query range. It highlights three different trends that for the low, medium, and high ranges are respectively the fastest, medium, and slowest convergence to the GT. To quantify the convergence speed to the GT, in Fig. 11b we plot the recall values at the point of the maximum convergence of the recall curves. In practice, it corresponds to the peak of the recall curve before its stability. If the maximum convergence rate is low ($<0.2$) for high recall values ($>0.8$), then it means that the approach converges slowly to the GT. It is the case of 100% of the experiments with high query ranges, 54% of the experiments with medium ranges, and 25% of the ones with low query ranges. In contrast, the rest of the medium and low-range queries show high convergence speed ($>0.3$) when the recall is higher than 0.8, indicating quick convergence to the GT. In Fig. 11c, we analyze the iterations remaining from peak convergence to retrieve all relevant paths (missing rate). The left part of the plot comprises experiments closer to recall value $1$, including all low-range experiments and 70% of medium-range ones. The right one involves cases needing more iterations to reach recall 1, encompassing all high-range experiments and the remaining medium-range ones. Results for low and medium ranges confirm SteerAG’s ability to quickly retrieve accurate attack paths even for strict queries. High-range query results seem inferior, but this is due to the large number of attack paths to retrieve, influenced by the sampling size. Precision remains consistently high, reaching convergence rates with recall values exceeding 0.8, even for high-range queries. Overall, we do not observe an effect on the performance.